OpenStudy (anonymous):

8/25 divide by 2 3/5

OpenStudy (anonymous):

Convert the second one to an improper fraction. Then flip it over and turn it in to a multiplication problem.

OpenStudy (anonymous):

so is that right

OpenStudy (anonymous):

So is what right?

OpenStudy (anonymous):

4/15

OpenStudy (anonymous):

Is the second number the mixed number "2 and 3 fifths"? Or is it "2 times 3 fifths"?

OpenStudy (anonymous):

mixed number

OpenStudy (anonymous):

Then 4/15 is not right. First you have to convert 2 3/5 to an improper fraction.

OpenStudy (anonymous):

so can u demonstrate it for me

OpenStudy (anonymous):

How about if you try it, and I'll help.

OpenStudy (anonymous):

2 3/5 would be 13/5

OpenStudy (anonymous):

Right. And to divide by a fraction, you just multiply by the reciprocal (upside-down version) of the fraction.

OpenStudy (anonymous):

multiply by 25/3

OpenStudy (anonymous):

Where did 25/3 come from?

OpenStudy (anonymous):

i mean 25/8

OpenStudy (anonymous):

No. Keep the first fraction the way it is. Turn the second one over.

OpenStudy (anonymous):

ok sooooo....8/25*5/13

OpenStudy (anonymous):

Exactly. Then do the multiplication and simplify the answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{8}{25} \div 2\frac{3}{5} = \frac{8}{25} \div \frac{13}{5} = \frac{8}{25} \times \frac{5}{13} = \frac{8}{65}\]

OpenStudy (anonymous):

okkk good thats what i got

OpenStudy (anonymous):

Good! (Now I know *I* didn't mess it up, either. ;) )